Shift summary for Owl, July 11, 1999

Angela Bellavance (bellavan@rice.edu)
Sun, 11 Jul 1999 07:00:46 -0500 (CDT)

Shift Summary for owl shift, 11/07/99

On Shift: Sasha Glasov, Angela Bellavance

Hours of beam delivered at all intensities: 6:07 (of 7)
Hours of beam delivered at nominal intensity: 6:07
Hours of E832 data (special runs not included): 5:58
Hours of special runs (laser scans, muon runs...): 0:11

Accounting of lost time:

Swap DPMT (from previous shift): 46 minutes
Problem with key tree: 7 minutes
Reset on spill scaler crate: 9 minutes

========================================================================
Runs ended this shift:
Run 13854 - PED
Run 13855 - 350k events - BEAM_832 (only ~8 good spills this run)
Run 13856 - still running - BEAM_832

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift:

At beginning, Hogan and Brad still on access to swap DPMT on channel 2575.

========================================================================
Overview of shift:
Mostly smooth E832 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
00:00 Hogan and Brad still on access to swap DPMT on
channel 2575.
13854 00:35 Do pedistal run to see that CsI is ok now and
load new constants.
13855 1 00:46 Start run 13855 but no beam yet. Problem with key
tree not closing ("A" LED off). People from MCR
to be sent out, but Sasha lays hands on the box
and fixes it. POOF! Yea!
13855 8 00:53 Beam returns, BOA:1.24
13855 15 01:04 DYC 32 full - kfrend fixes.
13855 ? 01:09 HCC ineff. warning. Reloading the Etot goes badly
and kfrend doesn't fix.
13855 ? 01:12 Stop run to reset crate 53 spill scaler.
Everything's ok now.
13856 1 01:18 Start run 13856.
13856 7 01:28 crate 53 full - kfrend fixes it.
02:20 Typing "yield832" in PAW gives N(VAC00)=1.03345K
02:36 Gas checklist done. It's 58 deg. outside! Brr...
07:00 Hand off to day shift.

========================================================================
Follow up Notes on Previous Shift:

Keep taking 832 data. No muon run needed.

========================================================================
Items to Follow up on Next Shift

TV monitor #413 for the beam position seems to be dying. A future day
shift needs to have it looked at.

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke